# Sequential Conditional Transport on Probabilistic Graphs

### v2

@@ Abstract
In this paper, we link two existing approaches to derive counterfactuals: adaptations based on a causal graph, and optimal transport. We extend “Knothe’s rearrangement" and “triangular transport" to probabilistic graphical models, and use this counterfactual approach, referred to as sequential transport, to discuss fairness at the individual level. After establishing the theoretical foundations of the proposed method, we demonstrate its application through numerical experiments on both synthetic and real datasets.
@@ 1 Introduction
Most applications concerning discrimination and fairness are based on “group fairness" concepts (as introduced in Hardt et al. (2016); Kearns and Roth (2019), or Barocas et al. (2023)). However, in many cases, fairness should be addressed at the individual level rather than globally. As claimed in Dwork et al. (2012), “we capture fairness by the principle that any two individuals who are similar with respect to a particular task should be classified similarly." The concept of “counterfactual fairness" was formalized in Kusner et al. (2017), addressing questions such as “had the protected attributes of the individual been different, other things being equal, would the decision had remain the same?" Such a statement has clear connections with causal inference, as discussed in Pearl and Mackenzie (2018). Formally, consider observations $\{s_{i},\boldsymbol{x}_{i},y_{i}\}$, where $s$ is a binary protected attribute (e.g., $s\in\{0,1\}$), and $\boldsymbol{x}$ is a collection of legitimate features (possibly correlated with $s$). The model output is $y$, which is analyzed to address “algorithmic fairness” issues. Following Rubin (2005), let $y^{\star}(s)$ denote the potential outcome of $y$ if $s$ is seen as a treatment. With these notations, counterfactual fairness is achieved for individual $(s,\boldsymbol{x})$ if the average “treatment effect,” conditional on $\boldsymbol{x}$ (or “CATE") is zero, i.e., $\mathbb{E}[Y^{\star}(1)-Y^{\star}(0)|\boldsymbol{X}=\boldsymbol{x}]=0$. This quantity could be termed “ceteris paribus CATE” since all $\boldsymbol{x}$’s are supposed to remain unchanged for both treated and non-treated.
@@
Following Kilbertus et al. (2017), it is possible to suppose that the protected attribute $s$ could actually affect some explanatory variables $\boldsymbol{x}$ in a non-discriminatory way. In Charpentier et al. (2023), the outcome $y$ was “having a surgical intervention" during childbirth in the U.S., $s$ was the mother’s ethnic origin (“Black," or not) and $\boldsymbol{x}$ included factors such as “weight of the baby at birth." If Black mothers undergo less surgery because they tend to have smaller babies, there is no discrimination per se. At the very least, it should be fair, when assessing whether hospitals have discriminatory policies, to account for that difference in baby weights. Such a variable is named “revolving variable" in Kilbertus et al. (2017). Using heuristic notations, the “ceteris paribus CATE" $\mathbb{E}[Y^{\star}(1)|\boldsymbol{X}=\boldsymbol{x}]-\mathbb{E}[Y^{\star}(0)|\boldsymbol{X}=\boldsymbol{x}]$ should become a “mutatis mutandis CATE". For some individual $(s=0,\boldsymbol{x})$, this indicator would be $\mathbb{E}[Y^{\star}(1)|\boldsymbol{X}=\boldsymbol{x}^{\star}(1)]-\mathbb{E}[Y^{\star}(0)|\boldsymbol{X}=\boldsymbol{x}]$, as coined in Charpentier et al. (2023), to quantify discrimination, where fictitious individual $(s=1,\boldsymbol{x}^{\star}(1))$ is a “counterfactual" version of $(s=0,\boldsymbol{x})$.
@@
Two recent approaches have been proposed to assess counterfactual fairness using this mutatis mutandis approach. On the one hand, Plečko and Meinshausen (2020) and Plečko et al. (2024) used causal graphs (DAGs) to construct counterfactuals and assess the counterfactual fairness of outcomes $y$ based on variables $(s,\boldsymbol{x},y)$. In network flow terminology, $s$ acts as a “source" (only outgoing flow, or no parents), while $y$ is a “sink" (only incoming flow). On the other hand, Black et al. (2020), Charpentier et al. (2023) and De Lara et al. (2024) used optimal transport (OT) to construct counterfactuals. Moreover, using counterfactual reasoning to achieve fair machine learning (ML) models has also been notably studied (Ma et al., 2023; Robertson et al., 2024). For evaluation, while De Lara et al. (2024) provided a theoretical framework, its implementation is challenging (except in the Gaussian case), and usually hard to interpret. Here, we combine the two approaches, using OT within a causal graph structure. The idea is to adapt “Knothe’s rearrangement" Bonnotte (2013), or “triangular transport” Zech and Marzouk (2022a, b), to a general probabilistic graphical model on $(s,\boldsymbol{x},y)$, rather than a simplistic $s\to x_{1}\to x_{2}\to\cdots\to x_{d}\to y$. The concept of “conditional OT" has been recently discussed in Bunne et al. (2022) and Hosseini et al. (2023), but here, instead of learning the causal graph, we assume a known causal graph and use it to construct counterfactual versions of individuals $(s_{i},\boldsymbol{x}_{i},y_{i})$ to address fairness issues. Additionally, since we use univariate (conditional) transport, standard classical properties of univariate transport facilitates explanations (non-decreasing mappings, and quantile based interpretations).
@@ 2.1 Probabilistic Graphical Models
Following standard notations in probabilistic graphical models (see Koller and Friedman (2009) or Barber (2012)), given a random vector $\boldsymbol{X}=(X_{1},\cdots,X_{d})$, consider a directed acyclic graph (DAG) $\mathcal{G}=(V,E)$, where $V=\{x_{1},x_{2},\cdots,x_{d}\}$ are the vertices (corresponding to each variable), and $E$ are directed edges, such that $x_{i}\to x_{j}$ means “variable $x_{i}$ causes variable $x_{j}$," in the sense of Susser (1991). The joint distribution of $\boldsymbol{X}$ satisfies the (global) Markov property w.r.t. $\mathcal{G}$:
@@
where $\text{parents}(x_{i})$ are nodes with edges directed towards $x_{i}$, in $\mathcal{G}$. Watson et al. (2021) suggested the causal graph in Figure 1 for the German Credit dataset, where $s$ is the “sex" (top left) and $y$ is the “default" indicator (right). Observe that variables $\boldsymbol{x}_{j}$ are here sorted. As discussed in Ahuja et al. (1993), such a causal graph imposes some ordering on variables. In this “topological sorting," a vertex must be selected before its adjacent vertices, which is feasible because each edge is directed such that no cycle exists in the graph. In our analysis, we consider a network $\mathcal{G}$ on variables $\{s,\boldsymbol{x},y\}$ where $s$ is the sensitive attribute, acting as a “source" (only outgoing flow, or no parents) while $y$ is a “sink" (only incoming flow, i.e., $y\notin\text{parents}(x_{i})$, $\forall i$).
@@ 2.2 Causal Networks and Linear Structural Models
Wright (1921, 1934) used directed graphs to represent probabilistic cause-and-effect relationships among a set of variables and developed path diagrams and path analysis. Simple causal networks can be visualized on top of Figure 2. On the left is a simple model where the “cause" $C$ directly causes ($\to$) the “effect" $E$. On the right, a “mediator" $X$ is added. There is still the direct impact of $C$ on $E$ ($C\to E$), but there is also a mediated indirect impact ($C\to X\to E$).
@@ 2.2.1 Intervention in a Linear Structural Model
where $U_{C}$ and $U_{E}$ are independent Gaussian random variables. That causal graph can be visualized in Figure 2, and its corresponding structural causal model (SCM) described in Equation 1 illustrates the causal relationships between variables, as in Pearl (2000). Suppose here that $C$ is a binary variable, taking values in $\{c_{0},c_{1}\}$. Given an observation $(c_{0},e)$, the “counterfactual outcome" if the cause had been set to $c_{1}$ (corresponding to the intervention in Figure 3), would be $e+b_{e}(c_{1}-c_{0})$. Following Pearl (2009), one can also introduce the “twin network" corresponding to a mirrored version of the initial causal graph in the counterfactual world. Plečko and Meinshausen (2020) coined this approach “fair-twin projection" when $C$ is a binary sensitive attribute.
@@ 2.3.2 Connections With Conditional Quantiles
This representation has been used in Plečko and Meinshausen (2020) and Plečko et al. (2024), where $X_{i}=F_{i|\boldsymbol{x}}^{-1}(V)$ is simply the probabilistic representation of “quantile regression," as introduced by Koenker and Bassett Jr (1978) (and further studied in Koenker (2005) and Koenker et al. (2017)). This could be extended to “quantile regression forests," as in Meinshausen and Ridgeway (2006), or any kind of ML model, as Cannon (2018) or Pearce et al. (2022). Observe that Ma and Koenker (2006) considered some close “recursive structural equation models," characterized by a system of equations where each endogenous variable is regressed on other endogenous and exogenous variables in a hierarchical manner. They used some sequential quantile regression approach to solve those recursive SEMs. An alternative we consider here is to use the connection between quantiles and OT (discussed in Chernozhukov et al. (2013) or Hallin and Konen (2024)) to define some “conditional transport" that relates to those conditional quantiles.
@@ 3 Optimal Transport
Out of those mappings from $\mu_{0}$ to $\mu_{1}$, we can be interested in “optimal" mappings, satisfying Monge problem, from Monge (1781), i.e., solutions of
@@ 4.1 Knothe-Rosenblatt Conditional Transport
As proved in Santambrogio (2015) and Carlier et al. (2010), $T_{\overline{kr}}$ is a transportation map from $\mu_{0}$ to $\mu_{1}$, in the sense that $\mu_{1}=T_{\overline{kr}\#}\mu_{0}$. Following Bogachev et al. (2005) and Backhoff et al. (2017), $T_{\overline{kr}}$ is the “monotone upper triangular map" uniquely defined when the $\mu_{1}$-conditionals are atomless for a chosen coordinate order. Bogachev et al. (2005) defined the “monotone lower triangular map,"
@@
The map $x_{i}\mapsto T^{\star}_{\underline{i}}(x_{i}|x_{1},\cdots,x_{i-1})$ is monotone (nondecreasing) for all $(x_{1},\cdots,x_{i-1})\in\mathbb{R}^{i-1}$. Further, by construction, this KR transport map has a triangular Jacobian matrix $\nabla T_{\underline{kr}}$ with nonnegative entries on its diagonal, making it suitable for various geometric applications. However, this mapping does not satisfy many properties; for example, it is not invariant under isometries of $\mathbb{R}^{d}$ as mentioned in Villani (2009). Carlier et al. (2010) proved that the KR transport maps could be seen as limits of quadratic OTs. A direct interpretation is that this iterative sequential transport can be seen as “marginally optimal." Some explicit formulas can be obtained in the Gaussian case, as discussed in Appendix A.
@@ 4.2 Sequential Conditional Transport on a Probabilistic Graph
The “monotone lower triangular map," introduced in Bogachev et al. (2005) could be used when dealing with time series, since there is a natural ordering between variables, indexed by the time, as discussed in Backhoff et al. (2017) or Bartl et al. (2021). In the general non-temporal case of time series $X_{t}$, it is natural to extend that approach to acyclical probabilistic graphic models, following Cheridito and Eckstein (2023). Instead of two general measures $\mu_{0}$ and $\mu_{1}$ on $\mathbb{R}^{d}$, we use only measures “factorized according to $\mathcal{G}$," some probabilistic graphical model, as defined in Lauritzen (2020).
@@
This mapping will be called “sequential conditional transport on the graph $\mathcal{G}$, "or shortly “sequential transport." [^1] Given the topological order of the graph and assuming the $\mu_{0},\mu_{1}$-conditionals are atomless, the existence and unicity of the sequential transport map are guaranteed, as it involves fewer conditioning variables compared to the KR transport map.
A classical algorithm for topological sorting is Kahn (1962)’s “Depth First Search" (DFS), and other algorithms are discussed in Section 20.4 in Cormen et al. (2022). For the causal graphs of Figure 5:
@@
In that simple case, for Figure 5a, we recognize the “monotone lower triangular map," and the “monotone upper triangular map," for 5b (see Section 4.1). Finally, for the causal graph on the German Credit dataset of Figure 1, variables are sorted, and
@@
Alternatively, using the “monotone lower triangular map" for the German Credit dataset to compute counterfactuals suggests that the assumed DAG contains more edges than the DAG illustrated in Figure 5. In this case, the edges are specified as $E=\{(i,j)\in V^{2}:i<j\}$, with $V=\{s,x_{1},x_{2},\dots,x_{7}\}$. Notably, multivariate OT corresponds to a fully connected graph, with $E=\{(i,j)\in V^{2}:i\neq j\}$ Cheridito and Eckstein (2023). The impact of edge mispecifications on sequential transport is examined in Appendix E. If the graph is entirely unknown, one could infer it as discussed in Zheng et al. (2018); Yu et al. (2019); Cai et al. (2023), or use Bayes’ rule to compute posterior DAGs, incorporating uncertainty quantification for counterfactuals Toth et al. (2022).

### v3

@@ Abstract
In this paper, we link two existing approaches to derive counterfactuals: adaptations based on a causal graph, and optimal transport. We extend “Knothe’s rearrangement” and “triangular transport” to probabilistic graphical models, and use this counterfactual approach, referred to as sequential transport, to discuss fairness at the individual level. After establishing the theoretical foundations of the proposed method, we demonstrate its application through numerical experiments on both synthetic and real datasets.
This works appears in the
Proceedings of the AAAI Conference on Artificial Intelligence, 39(18), 19358-19366.
doi: 10.1609/aaai.v39i18.34131
@@ 1 Introduction
Most applications concerning discrimination and fairness are based on “group fairness” concepts (as introduced in Hardt et al. (2016); Kearns and Roth (2019), or Barocas et al. (2023)). However, in many cases, fairness should be addressed at the individual level rather than globally. As claimed in Dwork et al. (2012), “we capture fairness by the principle that any two individuals who are similar with respect to a particular task should be classified similarly.” The concept of “counterfactual fairness” was formalized in Kusner et al. (2017), addressing questions such as “had the protected attributes of the individual been different, other things being equal, would the decision had remain the same?” Such a statement has clear connections with causal inference, as discussed in Pearl and Mackenzie (2018). Formally, consider observations $\{s_{i},\boldsymbol{x}_{i},y_{i}\}$, where $s$ is a binary protected attribute (e.g., $s\in\{0,1\}$), and $\boldsymbol{x}$ is a collection of legitimate features (possibly correlated with $s$). The model output is $y$, which is analyzed to address “algorithmic fairness” issues. Following Rubin (2005), let $y^{\star}(s)$ denote the potential outcome of $y$ if $s$ is seen as a treatment. With these notations, counterfactual fairness is achieved for individual $(s,\boldsymbol{x})$ if the average “treatment effect,” conditional on $\boldsymbol{x}$ (or “CATE”) is zero, i.e., $\mathbb{E}[Y^{\star}(1)-Y^{\star}(0)|\boldsymbol{X}=\boldsymbol{x}]=0$. This quantity could be termed “ceteris paribus CATE” since all $\boldsymbol{x}$’s are supposed to remain unchanged for both treated and non-treated.
@@
Following Kilbertus et al. (2017), it is possible to suppose that the protected attribute $s$ could actually affect some explanatory variables $\boldsymbol{x}$ in a non-discriminatory way. In Charpentier et al. (2023), the outcome $y$ was “having a surgical intervention” during childbirth in the U.S., $s$ was the mother’s ethnic origin (“Black” or not) and $\boldsymbol{x}$ included factors such as “weight of the baby at birth.” If Black mothers undergo less surgery because they tend to have smaller babies, there is no discrimination per se. At the very least, it should be fair, when assessing whether hospitals have discriminatory policies, to account for that difference in baby weights. Such a variable is named “resolving variable” in Kilbertus et al. (2017). Using heuristic notations, the “ceteris paribus CATE” $\mathbb{E}[Y^{\star}(1)|\boldsymbol{X}=\boldsymbol{x}]-\mathbb{E}[Y^{\star}(0)|\boldsymbol{X}=\boldsymbol{x}]$ should become a “mutatis mutandis CATE.” For some individual $(s=0,\boldsymbol{x})$, this indicator would be $\mathbb{E}[Y^{\star}(1)|\boldsymbol{X}=\boldsymbol{x}^{\star}(1)]-\mathbb{E}[Y^{\star}(0)|\boldsymbol{X}=\boldsymbol{x}]$, as coined in Charpentier et al. (2023), to quantify discrimination, where fictitious individual $(s=1,\boldsymbol{x}^{\star}(1))$ is a “counterfactual” version of $(s=0,\boldsymbol{x})$.
@@
Two recent approaches have been proposed to assess counterfactual fairness using this mutatis mutandis approach. On the one hand, Plečko and Meinshausen (2020) and Plečko et al. (2024) used causal graphs (DAGs) to construct counterfactuals and assess the counterfactual fairness of outcomes $y$ based on variables $(s,\boldsymbol{x},y)$. In network flow terminology, $s$ acts as a “source” (only outgoing flow, or no parents), while $y$ is a “sink” (only incoming flow). On the other hand, Black et al. (2020), Charpentier et al. (2023) and De Lara et al. (2024) used optimal transport (OT) to construct counterfactuals. Moreover, using counterfactual reasoning to achieve fair machine learning (ML) models has also been notably studied (Ma et al., 2023; Robertson et al., 2024). For evaluation, while De Lara et al. (2024) provided a theoretical framework, its implementation is challenging (except in the Gaussian case), and usually hard to interpret. Here, we combine the two approaches, using OT within a causal graph structure. The idea is to adapt “Knothe’s rearrangement” Bonnotte (2013), or “triangular transport” Zech and Marzouk (2022a, b), to a general probabilistic graphical model on $(s,\boldsymbol{x},y)$, rather than a simplistic $s\to x_{1}\to x_{2}\to\cdots\to x_{d}\to y$. The concept of “conditional OT” has been recently discussed in Bunne et al. (2022) and Hosseini et al. (2023), but here, instead of learning the causal graph, we assume a known causal graph and use it to construct counterfactual versions of individuals $(s_{i},\boldsymbol{x}_{i},y_{i})$ to address fairness issues. Additionally, since we use univariate (conditional) transport, standard classical properties of univariate transport facilitates explanations (non-decreasing mappings, and quantile based interpretations).
@@ 2.1 Probabilistic Graphical Models
Following standard notations in probabilistic graphical models (see Koller and Friedman (2009) or Barber (2012)), given a random vector $\boldsymbol{X}=(X_{1},\cdots,X_{d})$, consider a directed acyclic graph (DAG) $\mathcal{G}=(V,E)$, where $V=\{x_{1},x_{2},\cdots,x_{d}\}$ are the vertices (corresponding to each variable), and $E$ are directed edges, such that $x_{i}\to x_{j}$ means “variable $x_{i}$ causes variable $x_{j}$,” in the sense of Susser (1991). The joint distribution of $\boldsymbol{X}$ satisfies the (global) Markov property w.r.t. $\mathcal{G}$:
@@
where $\text{parents}(x_{i})$ are nodes with edges directed towards $x_{i}$, in $\mathcal{G}$. Watson et al. (2021) suggested the causal graph in Figure 1 for the German Credit dataset, where $s$ is the “sex” (top left) and $y$ is the “default” indicator (right). Observe that variables $\boldsymbol{x}_{j}$ are here sorted. As discussed in Ahuja et al. (1993), such a causal graph imposes some ordering on variables. In this “topological sorting,” a vertex must be selected before its adjacent vertices, which is feasible because each edge is directed such that no cycle exists in the graph. In our analysis, we consider a network $\mathcal{G}$ on variables $\{s,\boldsymbol{x},y\}$ where $s$ is the sensitive attribute, acting as a “source” (only outgoing flow, or no parents) while $y$ is a “sink” (only incoming flow, i.e., $y\notin\text{parents}(x_{i})$, $\forall i$).
@@ 2.2 Causal Networks and Linear Structural Models
Wright (1921, 1934) used directed graphs to represent probabilistic cause-and-effect relationships among a set of variables and developed path diagrams and path analysis. Simple causal networks can be visualized on top of Figure 2. On the left is a simple model where the “cause” $C$ directly causes ($\to$) the “effect” $E$. On the right, a “mediator” $X$ is added. There is still the direct impact of $C$ on $E$ ($C\to E$), but there is also a mediated indirect impact ($C\to X\to E$).
@@ 2.2.1 Intervention in a Linear Structural Model
where $U_{C}$ and $U_{E}$ are independent Gaussian random variables. That causal graph can be visualized in Figure 2, and its corresponding structural causal model (SCM) described in Equation 1 illustrates the causal relationships between variables, as in Pearl (2000). Suppose here that $C$ is a binary variable, taking values in $\{c_{0},c_{1}\}$. Given an observation $(c_{0},e)$, the “counterfactual outcome” if the cause had been set to $c_{1}$ (corresponding to the intervention in Figure 3), would be $e+b_{e}(c_{1}-c_{0})$. Following Pearl (2009), one can also introduce the “twin network” corresponding to a mirrored version of the initial causal graph in the counterfactual world. Plečko and Meinshausen (2020) coined this approach “fair-twin projection” when $C$ is a binary sensitive attribute.
@@ 2.3.2 Connections With Conditional Quantiles
This representation has been used in Plečko and Meinshausen (2020) and Plečko et al. (2024), where $X_{i}=F_{i|\boldsymbol{x}}^{-1}(V)$ is simply the probabilistic representation of “quantile regression,” as introduced by Koenker and Bassett Jr (1978) (and further studied in Koenker (2005) and Koenker et al. (2017)). This could be extended to “quantile regression forests,” as in Meinshausen and Ridgeway (2006), or any kind of ML model, as Cannon (2018) or Pearce et al. (2022). Observe that Ma and Koenker (2006) considered some close “recursive structural equation models,” characterized by a system of equations where each endogenous variable is regressed on other endogenous and exogenous variables in a hierarchical manner. They used some sequential quantile regression approach to solve those recursive SEMs. An alternative we consider here is to use the connection between quantiles and OT (discussed in Chernozhukov et al. (2013) or Hallin and Konen (2024)) to define some “conditional transport” that relates to those conditional quantiles.
@@ 3 Optimal Transport
Out of those mappings from $\mu_{0}$ to $\mu_{1}$, we can be interested in “optimal” mappings, satisfying Monge problem, from Monge (1781), i.e., solutions of
@@ 4.1 Knothe-Rosenblatt Conditional Transport
As proved in Santambrogio (2015) and Carlier et al. (2010), $T_{\overline{kr}}$ is a transportation map from $\mu_{0}$ to $\mu_{1}$, in the sense that $\mu_{1}=T_{\overline{kr}\#}\mu_{0}$. Following Bogachev et al. (2005) and Backhoff et al. (2017), $T_{\overline{kr}}$ is the “monotone upper triangular map” uniquely defined when the $\mu_{1}$-conditionals are atomless for a chosen coordinate order. Bogachev et al. (2005) defined the “monotone lower triangular map,”
@@
The map $x_{i}\mapsto T^{\star}_{\underline{i}}(x_{i}|x_{1},\cdots,x_{i-1})$ is monotone (nondecreasing) for all $(x_{1},\cdots,x_{i-1})\in\mathbb{R}^{i-1}$. Further, by construction, this KR transport map has a triangular Jacobian matrix $\nabla T_{\underline{kr}}$ with nonnegative entries on its diagonal, making it suitable for various geometric applications. However, this mapping does not satisfy many properties; for example, it is not invariant under isometries of $\mathbb{R}^{d}$ as mentioned in Villani (2009). Carlier et al. (2010) proved that the KR transport maps could be seen as limits of quadratic OTs. A direct interpretation is that this iterative sequential transport can be seen as “marginally optimal.” Some explicit formulas can be obtained in the Gaussian case, as discussed in Appendix A.
@@ 4.2 Sequential Conditional Transport on a Probabilistic Graph
The “monotone lower triangular map,” introduced in Bogachev et al. (2005) could be used when dealing with time series, since there is a natural ordering between variables, indexed by the time, as discussed in Backhoff et al. (2017) or Bartl et al. (2021). In the general non-temporal case of time series $X_{t}$, it is natural to extend that approach to acyclical probabilistic graphic models, following Cheridito and Eckstein (2023). Instead of two general measures $\mu_{0}$ and $\mu_{1}$ on $\mathbb{R}^{d}$, we use only measures “factorized according to $\mathcal{G}$,” some probabilistic graphical model, as defined in Lauritzen (2020).
@@
This mapping will be called “sequential conditional transport on the graph $\mathcal{G}$,” or shortly “sequential transport.” [^1] Given the topological order of the graph and assuming the $\mu_{0},\mu_{1}$-conditionals are atomless, the existence and unicity of the sequential transport map are guaranteed, as it involves fewer conditioning variables compared to the KR transport map.
A classical algorithm for topological sorting is Kahn (1962)’s “Depth First Search” (DFS), and other algorithms are discussed in Section 20.4 in Cormen et al. (2022). For the causal graphs of Figure 5:
@@
In that simple case, for Figure 5a, we recognize the “monotone lower triangular map,” and the “monotone upper triangular map,” for 5b (see Section 4.1). Finally, for the causal graph on the German Credit dataset of Figure 1, variables are sorted, and
@@
Alternatively, using the “monotone lower triangular map” for the German Credit dataset to compute counterfactuals suggests that the assumed DAG contains more edges than the DAG illustrated in Figure 5. In this case, the edges are specified as $E=\{(i,j)\in V^{2}:i<j\}$, with $V=\{s,x_{1},x_{2},\dots,x_{7}\}$. Notably, multivariate OT corresponds to a fully connected graph, with $E=\{(i,j)\in V^{2}:i\neq j\}$ Cheridito and Eckstein (2023). The impact of edge mispecifications on sequential transport is examined in Appendix E. If the graph is entirely unknown, one could infer it as discussed in Zheng et al. (2018); Yu et al. (2019); Cai et al. (2023), or use Bayes’ rule to compute posterior DAGs, incorporating uncertainty quantification for counterfactuals Toth et al. (2022).